it's been awhile since I've worked out story problems, my step daughter came home today with some and I'm hoping I could get some help refreshing my memory how they are done,so I can help her.I vaguely remember using a box to solve, any how here is the first one.
Admission to a baseball game is $3.00 for general admission and $6.00 for reserved seats. The receipts were $3561.00 for 868 paid admissions. how many of each ticket were sold?

Answer by ankor@dixie-net.com(22740)

Admission to a baseball game is $3.00 for general admission and $6.00 for reserved seats.
The receipts were $3561.00 for 868 paid admissions.
how many of each ticket were sold?
:
The easiest way to solve this is with a system of equations
:
Let g = no. of general admission tickets
Let r = no. of reserve tickets
:
The total admissions equation
g + r = 868
g = (868-r); use this form for substitution
:
The total receipts equation
3g + 6r = 3561
substitute g with (868-r)
3(868-r) + 6r = 3561
2604 - 3r + 6r = 3561
-3r + 6r = 3561 - 2604
3r = 957
r = 957/3
r = 319 reserve seat
then
868 - 319 = 549 general admission seats
:
:
Confirm this in the total receipts equation